OpenStudy (anonymous):

find 4 arithmetic means between 2 and 37

OpenStudy (anonymous):

Let a1,a2,a3,a4 be 4 A.M.S then 2,a1,a2,a3,a4,37 are in A.P. Let d be c.d. here no. of terms n=6 first term a=2 now tn=a+(n-1)d 37=2+(6-1)d d=35/5=7 each A.M=a+nd,substitute a=2,d=7,n=1,2,3,4 we get the reqd. nos.